OpenStudy (anonymous):

A square is inscribed in a circle. If the radius of the circle is increasing at a rate of 3 cm/min, find how fast the area of the square is increasing when the radius is 10 cm.

OpenStudy (rulnick):

OK, here goes:

OpenStudy (rulnick):

r=d/2 s=d/sqrt(2) a=s^2=2r^2 da/dr = 4r da/dt = da/dr dr/dt = 4r dr/dt = 4r (3 cm/min) = 12r cm/min at r=10cm this is da/dt = 120 cm^2/min

OpenStudy (anonymous):

thank you =)

OpenStudy (rulnick):

You're welcome. Good problem. Points for your math teacher.